OpenStudy (anonymous):

Does anyone know the integral of cos(2x)?

OpenStudy (anonymous):

yes, it's like chain rule but you divide instead of multiplying

OpenStudy (anonymous):

1/2 sin(2x)

OpenStudy (anonymous):

:O Thanksomuch

OpenStudy (across):

You could also derive it by substituting:\[\int\cos(2x)dx,\]\(u=2x\), \(du=2dx\),\[\frac{1}{2}\int\cos(u)du=\frac{1}{2}\sin(2x).\]

OpenStudy (across):

... + c ^^